1.David's bokstore foresees that people will buy notebooks several days before the start of a new school year. The cost of the notebook in the store is P25. Two weeks before the start of the new school year, the store's rate of markup based on the selling price is 8%. One week after the opening of classes, they decided to put additional mark-on of 2% to its selling price.



A. What is the selling price?


B. What is the markup?


C. What is the additional mark-on?


D. How much is the new selling price after the additional mark-on?

1
Expert's answer
2021-11-12T04:01:45-0500

Solution :- markup % =


[(selling price - cost price)/cost price ]×100


this implies 8 = [(SP - 25)/25]× 100


8× 25 = (SP -25) × 100


8×25/100 = SP -25


2 = SP -25


2+25 = SP


so, SP = Php 27


(2) Markup = = 27-25 = 2


(3) The additional mark on = 2% of 27


= 0.54


(4) The new selling price = 27+0.54


= 27.54